# DiagGPT: An LLM-based and Multi-agent Dialogue System with Automatic Topic Management for Flexible Task-Oriented Dialogue

### v3

@@ Appendix B Analysis of DiagGPT Extendibility
We have only extracted the most important modules in DiagGPT to form a basic framework of a system. These four modules can already implement the basic functions of task-oriented dialogue. In a multi-agent AI system, there is a large extendibility. For example, an information collector can monitor user input and organize information into structured data for better future utilization.
@@
Sometimes, a conversation is aimed at solving certain tasks during or after the conversation. When a conversation achieves the predefined goal, the system can call more complex programs to meet needs. Some tool API (Application Programming Interface) calls can also be added into the action list for execution, which means the action list is the interface of DiagGPT and can provide many plugins to enrich the functions of this AI system. This also means that it can expand task-oriented dialogue to task-oriented dialogue by triggering the API of certain tasks.
Figure 6: A dialogue example in the medical consulting scenario. The AI system acts as a real doctor in our daily life.
Figure 8: A dialogue example in the medical consulting scenario when user ask some questions about COVID-19.
Appendix C Implementation Details of DiagGPT
@@
In the implementation of DiagGPT, we employed gpt-4 as the base LLM, leveraging its strong understanding and reasoning abilities to achieve ideal results. We set the decoding temperature of all LLMs in our AI system to 0 to ensure more stable task execution. We provide detailed prompts in our AI system. The main prompts of the AI agent are shown in Figure 10, while Figure 11 displays the prompts for all actions in the action list. In these prompts, {variable} in blue indicates that the slot needs to be filled with the corresponding variable text. Once filled, these prompts can be fed into the LLMs to generate responses. Some of these slots facilitate information interaction between AI systems.
Appendix D Qualitative Experiment Results
Figure 6 and Figure 7 present a complete dialogue demonstration in the medical consulting process. This is a medical diagnosis where the task is to help the patient identify the cause and give advice. Medical consulting is not a pure open-end question-answering. Because patients usually lack extensive medical knowledge, they rely on doctors to instruct them to give their personal information. Therefore, users can experience a real doctor, not a dull medical question-answering machine. The user acts as a patient, while the AI emulates a doctor, initially collecting information and gradually providing advice to the patient. The main dialogue development follows the checklist: Basic information, Chief complaint, Duration of symptoms, Severity of symptoms, which are also predefined topics.
This demonstration has showcased the robust conversational ability of the DiagGPT, which can actively ask questions and guide the user to the final goal of the task, thereby achieving task-oriented dialogue. It simulates many real consulting scenarios. Other chat models, such as ChatGPT, cannot achieve this performance. They usually just answer user questions and find it challenging to complete specific goals, even with elaborated prompts.
Appendix E Case Study of Automatic Topic Management
The core capability of DiagGPT is to automatically manage topics throughout the dialogue, which is the function and primary responsibility of Topic Manager. Figure 6 and Figure 7 illustrate the primary checklist progression. The topics from the checklist are retrieved and discussed sequentially, demonstrating the action of finishing the current topic. When the conversation reaches the severity of symptoms, we observe that the dialogue topic remains here in several rounds of dialogue, allowing the system to have time on understanding the user’s conditions. This mirrors real-world scenarios where users do not provide enough information for the doctor. The action of creating a new topic is shown in Figure 8. Here, the user actively consults the system with some information about COVID-19 to check symptoms. We observe that the AI generates a new topic about COVID-19 and discusses this, rather than rigidly following the checklist. These results all demonstrate the effectiveness of Topic Manager. These case studies fully embody the AI’s flexible understanding ability, demonstrating its adept handling of different situations, closely mirroring real-world interactions.
Appendix F LLM-TOD Dataset
@@
We have constructed a new dataset, LLM-TOD, to evaluate the performance of DiagGPT qualitatively. Since DiagGPT relies on LLMs, and considering our limited resources, fine-tuning LLMs for this specific purpose of TOD is not feasible. Therefore, we are unable to utilize other datasets that are tailored for fine-tuning models. The dataset comprises 20 data, each representing a different topic: clinical, restaurant, hotel, hospital, train, police, bus, attraction, airport, bar, library, museum, park, gym, cinema, office, barbershop, bakery, zoo, and bank. Each data entry is generated by LLMs and subsequently validated by humans. For every task-oriented dialogue in the dataset, the system is required to go through a checklist in order and interact with users to achieve the specified goal. We incorporate these variables into the prompt for LLMs, thereby enabling them to execute the TOD. An example from the TOD-LLM dataset is illustrated in Figure 9.
Appendix G Prompt Design
The content of prompts for some LLMs in DiagGPT are shown in Figure 10 and Figure 11.
Figure 10: The prompts of Chat Agent, Topic Manager, Topic Enricher. We have included instructions to guide the AI in becoming a knowledgeable medical expert, making it applicable in medical dialogue scenarios. These instructions can be modified to suit other scenarios.

### v4

@@ Appendix A Analysis of DiagGPT Extendibility
In this paper, we only introduce the basic framework of the system DiagGPT aimed at achieving task-oriented dialogue. We have designed the system with ample flexibility to incorporate additional functions to handle tasks in complex scenarios and to meet more needs of conversational systems. Besides, as the foundation model is upgraded, the performance of our system will also become better. In detail, we have only extracted the most important modules in DiagGPT to form a basic framework of a system. These four modules can already implement the basic functions of task-oriented dialogue. In a multi-agent system, there is a large extendibility. For example, an information collector can monitor user input and organize information into structured data for better future utilization.
@@
Sometimes, a conversation is aimed at solving certain tasks during or after the conversation. When a conversation achieves the predefined goal, the system can call more complex programs to meet needs. Some tool API (Application Programming Interface) calls can also be added into the action list for execution, which means the action list is the interface of DiagGPT and can provide many plugins to enrich the functions of this system. This also means that it can expand task-oriented dialogue to task-oriented dialogue by triggering the API of certain tasks.
Appendix B Implementation Details of DiagGPT
@@
In the implementation of DiagGPT, we employed gpt-4 as the base LLM, leveraging its strong understanding and reasoning abilities to achieve ideal results. We set the decoding temperature of all LLMs in our AI system to 0 to ensure more stable task execution. We provide detailed prompts in our system. The main prompts of the agent are shown in Figure 10, while Figure 11 displays the prompts for all actions in the action list. In these prompts, {variable} in blue indicates that the slot needs to be filled with the corresponding variable text. Once filled, these prompts can be fed into the LLMs to generate responses. Some of these slots facilitate information interaction between agents in the system.
In the evaluation, we use gpt-4-0613 as the foundation model for DiagGPT and for ChatGPT (gpt-4). For ChatGPT (gpt-3.5-turbo), we select gpt-3.5-turbo-0613. The simulation of the user is also based on gpt-4-0613. For the LLM evaluation, we use gpt-4-turbo-2024-04-09. To simplify the experimental setup and facilitate easier comparisons, we use a simplified version of DiagGPT in the quantitative experiment. We reduce the complexity of the action list and optimized prompts of LLM to promote topic development as much as possible. However, in qualitative experiments, we evaluate the full version of DiagGPT, which showcases the complete capabilities of our system.
Appendix C LLM-TOD Dataset
@@
We construct a new dataset, LLM-TOD (task-oriented dialogue for large language models dataset). It is used to evaluate the performance of LLM-based task-oriented dialogue models quantitatively. Since DiagGPT relies on LLMs, and considering our limited resources, fine-tuning LLMs for this specific purpose of TOD is not feasible. Therefore, we are unable to utilize other datasets that are tailored for fine-tuning models.
The dataset comprises 20 data, each representing a different topic: clinical, restaurant, hotel, hospital, train, police, bus, attraction, airport, bar, library, museum, park, gym, cinema, office, barbershop, bakery, zoo, and bank. Each data entry is generated by LLMs and subsequently validated by humans. For every task-oriented dialogue in the dataset, the system is required to go through a checklist in order and interact with users to achieve the specified goal. One checklist includes six items of topic or task and need to be solved or discussed. We incorporate these variables into the prompt for LLMs, thereby enabling them to execute the TOD. An example from the TOD-LLM dataset is illustrated in Figure 9.
Appendix D Prompt Design
The content of prompts for some LLMs in DiagGPT are shown in Figure 10 and Figure 11. Figure 12 also displays the prompts used for other LLM-based models in our experiments: the ChatGPT baselines, UserGPT (the user simulator), and EvalGPT (used for LLM evaluation). Grading mode involves assigning a score to the quality of responses, while comparison mode requires selecting a winner based on response quality in a comparison. The criteria for evaluating response quality from both human and LLM perspectives include:
  • Understanding How well does the system grasp user requests?
  • Relevance Are the responses directly applicable to the user’s needs?
  • Complex Handling Can the system effectively manage multifaceted queries?
  • Efficiency How quickly does the system lead the conversation to a resolution?
  • Experience What is the overall ease and satisfaction of the interaction?
  • Comprehensiveness Does the system’s response provide all the relevant information to fully cover the user’s question?
  • Detail Does the response include sufficient detail and depth to fully address the intricacies and subtleties of the user’s inquiry?
  • Sufficiency Are all aspects and potential implications of the inquiry explored and explained, ensuring that the user receives a thorough understanding of the subject matter?
Figure 10: The prompts of Chat Agent, Topic Manager, Topic Enricher. These instructions can be modified to suit other scenarios.